# EgoTV [Uncaptioned image]: Egocentric Task Verification from Natural Language Task Descriptions

### v5

@@ 2 Related Work
Neuro-symbolic Models. Neuro-symbolic models combine feature extraction through deep learning with symbolic reasoning [nesy-survey, star_situated_reasoning] to capture compositional substructures. These models either reason on static images to recognize object attributes and relations (NS-CL [nscl], NS-VQA [nsvqa], CLOSURE [closure], and $\nabla$-∇ -FOL [del-fol]), or on videos to recognize spatio-temporal and causal relations (NS-DR [clevrer] and DCL [dcl]). We extend this to tracking multi-step actions.
@@ 3.1 Definitions
[…]
@@ 5.3 Query Encoders for Grounding
Figure 4: NSG model (a) semantic parser converts NL descriptions into a graph $G$ of symbolic queries; query encoders $f^{\theta_{\tau}}$ detect queries in individual video segments $s_{t}$; and a video aligner aligns $G$ with video segments by computing alignment matrix $\mathrm{Z}$ via a constrained optimization problem (Eq. 5.4). (b) The constraints (Eqs. 3a 3b 3c) and the recursive structure (Eq. 4) enabling use of DP to solve for $\mathrm{Z}$. Here, the blue box denotes $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}}^{S-1})$, the green boxes denote $\log f^{\theta}(a_{\bar{j}},s_{\bar{t}})+F^{\ast}((n_{j})_{\bar{j}+1}^{N-1},(s_{t})_{\bar{t}+1}^{S-1})$, and the red box denotes $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}+1}^{S-1})$.
@@ 5.4 Video Aligner for Task Verification
[…]
Video Segmentation: The video is segmented into non-overlapping segments [^3] Since pretrained, off-the-shelf video segmentation models are limited to predefined action classes [escorcia2016daps] or reliant on background frame change detection [yang2022temporal] and require downstream finetuning [gao2020accurate], we leave their integration in NSG as future work. with a moving window of arbitrary, but fixed size $k$ [^4] If required, the last segment is zero-padded to $k$ frames.
@@
[…]
@@
Intuitively, the solution to Eq. 5.4 gives us the best alignment score (note, the overlap with Eq. 1). The iterations over $N$ queries and $S$ segments for solving Eq. 5.4 are underpinned by an overlapping and optimal substructure. For instance, to optimally align queries $(n_{j})_{j=0}^{N-1}$ and segments $(s_{t})_{t=0}^{S-1}$, one could: a) pair $\langle n_{0},s_{0}\rangle$ and optimally align the remaining queries and segments $(n_{j})_{j=1}^{N-1},(s_{t})_{t=1}^{S-1}$; or (2) skip $s_{0}$ and still optimally align all queries, now with the remaining segments $(n_{j})_{j=0}^{N-1},(s_{t})_{t=1}^{S-1}$ (see Fig. 4b(iv)). This recursive substructure leads to a DP solution for Eq. 5.4.
Let, $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}}^{S-1})$ denote the best alignment score for queries $(n_{j})_{\bar{j}}^{N-1}$ and segments $(s_{t})_{\bar{t}}^{S-1}$ from Eq. 5.4. Based on the aforementioned reasoning, $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}}^{S-1})$ can be recursively written as:

### v6

@@ 2 Related Work
Neuro-symbolic Models. Neuro-symbolic models combine feature extraction through deep learning with symbolic reasoning [nesy-survey, star_situated_reasoning] to capture compositional substructures. These models either reason on static images to recognize object attributes and relations (NS-CL [nscl], NS-VQA [nsvqa], CLOSURE [closure], and $\nabla$-FOL [del-fol]), or on videos to recognize spatio-temporal and causal relations (NS-DR [clevrer] and DCL [dcl]). We extend this to tracking multi-step actions.
@@ 3.1 Definitions
The ordering constraints between sub-tasks are captured in the task description using specifiers such as and, then, and before/after. For simplicity, we will refer to a task using $\left<\text{\emph{sub-task}}\right>\_\left<\text{\emph{ordering-specifier}}\right$> notation, irrespective of the actual task description. Such tasks can then be instantiated by specifying an $(object)$ of interaction. An example task instance from EgoTV:  heat_then_clean(apple) is shown in Fig. 1 with its NL description: “apple is heated, then cleaned in a sinkbasin”. The task consists of two ordered sub-tasks: heat $\rightarrow$ clean on target object: apple. We adopt this terminology from ALFRED [ALFRED20].
@@ 5.3 Query Encoders for Grounding
Figure 4: NSG model (a) semantic parser converts NL descriptions into a graph $G$ of symbolic queries; query encoders $f^{\theta_{\tau}}$ detect queries in individual video segments $s_{t}$; and a video aligner aligns $G$ with video segments by computing alignment matrix $\mathrm{Z}$ via a constrained optimization problem (Eq. 3). (b) The constraints (Eqs. 3a 3b 3c) and the recursive structure (Eq. 4) enabling use of DP to solve for $\mathrm{Z}$. Here, the blue box denotes $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}}^{S-1})$, the green boxes denote $\log f^{\theta}(a_{\bar{j}},s_{\bar{t}})+F^{\ast}((n_{j})_{\bar{j}+1}^{N-1},(s_{t})_{\bar{t}+1}^{S-1})$, and the red box denotes $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}+1}^{S-1})$.
@@ 5.4 Video Aligner for Task Verification
This module of NSG must align the graph representation $G$ of the task (generated by the semantic parser) with the video. To that end, it first segments the video, then jointly learns a) the query encoders, which detect the queries in the video segments and b) the alignment between video segments and the query sequences obtained from the topological sort on $G$. Such joint learning is required since the temporal locations of the queries in the video are unknown a priori requiring simultaneous detection and alignment. If the video is a positive match for the task encoded in $G$, at least one of the query sequences from $G$ must temporally align perfectly with the video segments for successful task verification. Conversely, for negative matches, no query sequence from $G$ would completely align with the video segments. Going forward, we use $\langle\rangle$ and $()$ to denote ordered pairs and sequences, respectively.
Video Segmentation: The video is segmented into non-overlapping segments [^3] Since pretrained, off-the-shelf video segmentation models are limited to predefined action classes [escorcia2016daps] or reliant on background frame change detection [yang2022temporal] and require downstream finetuning [gao2020accurate], we leave their integration in NSG as future work. with a moving window of arbitrary, but fixed size kk [^4] If required, the last segment is zero-padded to $k$ frames.
@@
$\max_{\mathrm{Z}\in{\{0,1\}}^{N\times S}}\sum_{j,t}\log f^{\theta}(a_{j},s_{t})Z_{jt},\quad\text{s.t.}$ (3)
$\sum_{j=0}^{N-1}Z_{jt}\in\{0,1\},\quad\forall\;0\leq t\leq S-1$ (3a)
$\sum_{t=0}^{S-1}Z_{jt}=1,\quad\forall\;0\leq j\leq N-1$ (3b)
@@
Intuitively, the solution to Eq. 3 gives us the best alignment score (note, the overlap with Eq. 1). The iterations over $N$ queries and $S$ segments for solving Eq. 3 are underpinned by an overlapping and optimal substructure. For instance, to optimally align queries $(n_{j})_{j=0}^{N-1}$ and segments $(s_{t})_{t=0}^{S-1}$, one could: a) pair $\langle n_{0},s_{0}\rangle$ and optimally align the remaining queries and segments $(n_{j})_{j=1}^{N-1},(s_{t})_{t=1}^{S-1}$; or (2) skip $s_{0}$ and still optimally align all queries, now with the remaining segments $(n_{j})_{j=0}^{N-1},(s_{t})_{t=1}^{S-1}$ (see Fig. 4b(iv)). This recursive substructure leads to a DP solution for Eq. 3.
Let, $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}}^{S-1})$ denote the best alignment score for queries $(n_{j})_{\bar{j}}^{N-1}$ and segments $(s_{t})_{\bar{t}}^{S-1}$ from Eq. 3. Based on the aforementioned reasoning, $F^{\ast}((n_{j})_{\bar{j}}^{N-1},(s_{t})_{\bar{t}}^{S-1})$ can be recursively written as: